# Quantization-free Lossy Image Compression

### v1

@@ 5.  Discussion
All our comparative results (Figure 3 and Figure 5) consistently show that our IMF method outperforms JPEG in both maintaining image quality and preserving visual semantics at low bit rates and remains comparable at higher bit rates. Moreover, IMF consistently demonstrates superior performance compared to SVD across all bit rates. This superiority can be attributed to the quantization-free nature of IMF, which allows for more accurate reconstruction. In contrast, the high sensitivity of SVD to quantization errors during encoding and decoding degrades reconstruction quality.
@@
As observed in Figure 6(a), contracting the IMF factor bounds from $[-128,127]$ to $[-16,15]$ consistently improves the rate-distortion performance. Generally, narrowing the factor bounds $[\alpha,\beta]$ can potentially lower the entropy, thereby improving the effectiveness of lossless compression in the final stage of our framework and subsequently reducing the bit rate. However, this reduction in entropy comes at the cost of increased reconstruction error, as the feasible set in (3.3) becomes more constrained. This trade-off between entropy and reconstruction quality limits the compression performance of IMF. Therefore, it would be beneficial to moderately expand the factor bounds $[\alpha,\beta]$ while simultaneously controlling the entropy of the elements in the factor matrices. We plan to address this in the future by incorporating an entropy-aware regularization term into the current IMF objective function.
@@
Patchification with an appropriate patch size (e.g., $(8,8)$) helps capture local spatial dependencies and, as confirmed by our results in Figure 6(c), positively impacts the performance of IMF and SVD. However, discontinuities at patch boundaries can introduce blocking artifacts, similar to JPEG compression at very low bit rates (see the building image example in Figure 4). Moreover, while JPEG suffers more from color distortion (e.g., color bleeding and color banding) at low bit rates, IMF and SVD are more affected by blurriness, as observed in the seascape image example in Figure 4. As a potential solution for future work, a deep neural network could be trained to remove these artifacts and then integrated as a post-processing module to further enhance the quality of IMF-compressed images.
6.  Conclusion
@@
This work presents a novel quantization-free lossy image compression method based on integer matrix factorization (IMF). By representing image data as a product of two smaller matrices with bounded integer elements, the proposed IMF approach effectively eliminates the need for quantization, a significant source of error in traditional compression methods like JPEG and SVD. The reshaped factor matrices are compatible with existing lossless compression standards, enhancing the overall flexibility and efficiency of our method. Our proposed iterative algorithm, utilizing a block coordinate descent scheme, has proven to be both efficient and convergent. Experimental results demonstrate that the IMF method significantly outperforms JPEG in terms of PSNR and SIMM at low bit rates and maintains better visual semantic information. This advantage underscores the potential of IMF to set a new standard in lossy image compression, bridging the gap between factorization and quantization.
Appendix A Proof of Theorem 3.1
@@
$\operatorname*{arg\,min}_{u^{r}_{i}\in\mathbb{Z}_{[\alpha,\beta]}}$ $~{}q_{i}(u_{i}^{r}),\quad\forall i\in\{1,\dots,M\},$
where $q_{i}(u_{i}^{r})\coloneqq\sum_{j=1}^{N}(e_{ij}^{r}-u_{i}^{r}v_{j}^{r})^{2}.$ (11)
@@
Since $\bm{u}_{r}^{\star}\coloneqq(u^{r^{\star}}_{1},\dots,u^{r^{\star}}_{M})$ is the global optimum of optimization (10), it is evident that
@@
Appendix B Proof of Theorem 3.2
@@
$\Psi(\bm{U},\bm{V})$ $\coloneqq f_{0}(\bm{U},\bm{V})+\sum_{r=1}^{R}f(U_{:r})+\sum_{r=1}^{R}g(V_{:r}),$
$f_{0}(\bm{U},\bm{V})$ $\coloneqq\|\bm{X}-\bm{U}\bm{V}^{\rm T}\|_{\rm F}^{2},$
$f(U_{:r})$ $\coloneqq\delta_{[a,b]}(U_{:r})+\delta_{\mathbb{Z}}(U_{:r}),$
$g(V_{:r})$ $\coloneqq\delta_{[a,b]}(V_{:r})+\delta_{\mathbb{Z}}(V_{:r}),$
@@
The unconstrained optimization problem (14) consists of the sum of a differentiable (smooth) function $f_{0}$ and nonsmooth functions $f$ and $g$. This problem has been extensively studied in the literature under the class of nonconvex nonsmooth minimization problems. In Algorithm 1, the blocks $U_{:r}$ and $V_{:r}$ are updated sequentially following block coordinate descent (BCD) minimization algorithms, also often called Gauss-Seidel updates or alternating optimization [nesterov2012efficiency, attouch2013convergence]. Hence, in this convergence study, we are interested in algorithms that allow BCD-like updates for the nonconvex nonsmooth problem of (14) [beck2013convergence, bolte2014proximal]. Specifically, we focus on the proximal alternating linearized minimization (PALM) algorithm [bolte2014proximal], to relate its convergence behavior to that of Algorithm 1. To that end, we show that the updates of Algorithm 1 are related to the updates of PALM on the recast problem of (14), and all the assumptions necessary for the convergence of PALM are satisfied by our problem setting. It is noted that, for the sake of presentation and without loss of generality, in this proof, we assume each of the matrices $\bm{U}$ and $\bm{V}$ has only one column ($R=1$); hence, we only have two blocks in the BCD updates. The iterates in PALM and the presented proof can be trivially extended for more than two blocks.
@@
where the proximal map for an extended proper lower semicontinuous (nonsmooth) function $\varphi:\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ and $\gamma>0$ is defined as $\operatorname{prox}^{\varphi}_{\gamma}(\bm{x})\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\mathbb{R}^{n}}\left\{\varphi(\bm{w})+\frac{\gamma}{2}\|\bm{w}-\bm{x}\|^{2}_{2}\right\}$. In (15), $c_{k}>L_{1}(\bm{V}^{k})$ and $d_{k}>L_{2}(\bm{U}^{k+1})$ where $L_{1}>0$, $L_{2}>0$ are local Lipschitz moduli, defined in the following proposition.
The following proposition investigates the necessary assumptions (cf. [bolte2014proximal, Asm. 1 and Asm. 2]) for convergence of iterates in (15).
Proposition B.1 (Meeting required assumptions).
@@
    $\|\big{(}\nabla_{\bm{U}}f_{0}(\bm{U}_{1},\bm{V}_{1})-\nabla_{\bm{U}}$ $f_{0}(\bm{U}_{2},\bm{V}_{2}),$
    $\nabla_{\bm{V}}f_{0}(\bm{U}_{1},\bm{V}_{1})$ $-\nabla_{\bm{V}}f_{0}(\bm{U}_{2},\bm{V}_{2})\big{)}\|$
    $\leq M\|(\bm{U}_{1}-\bm{U}_{2},\bm{V}_{1}-\bm{V}_{2})\|,$
@@
By Proposition B.1, the optimization problem (14) can be solved by the iterates in (15), due to the following proposition:
Proposition B.2 (Global convergence [bolte2014proximal]).
With the assumptions in proposition B.1 being met by the problem (14), let $((\bm{U}^{k},\bm{V}^{k}))_{k\in\mathbb{N}}$ be a sequence generated by the iterates in (15). Then the sequence converges to a critical point $(\bm{U}^{\star},\bm{V}^{\star})$ of the problem (14), where $0\in\partial\Psi(\bm{U}^{\star},\bm{V}^{\star})$, with $\partial$ as the subdifferential of $\Psi$.
@@
Lemma B.1 ($\operatorname{prox}$ implementation).
@@
$\|\bm{W}\|_{[a,b]}^{2}\coloneqq\sum_{i,j\mid a\leq\bm{W}_{ij}\leq b}\bm{W}_{ij}^{2},\quad\|\bm{W}\|_{a}^{2}\coloneqq\sum_{i,j\mid\bm{W}_{ij}<a}\bm{W}_{ij}^{2},$
$\|\bm{W}\|_{b}^{2}\coloneqq\sum_{i,j\mid\bm{W}_{ij}>b}\bm{W}_{ij}^{2}.$
@@
Now that the equivalence of iterates (15) with the simple and closed-form steps in Algorithm 1 is fully established, and the assumptions required for the convergence are verified in proposition B.1 to be met by problems (14) and (3.3), proposition B.2 can be trivially invoked to establish the convergence of Algorithm 1 to a locally optimal point of problem (3.3).
[…]

### v2

@@ 5 Discussion
All our comparative results (Figure 3 and Figure 5) consistently show that our QMF method outperforms JPEG in both maintaining image quality and preserving visual semantics at low bit rates and remains comparable at higher bit rates. Moreover, QMF consistently demonstrates superior performance compared to SVD across all bit rates. This superiority can be attributed to the integration of quantization with low-rank approximation in QMF, which enables more accurate reconstruction. In contrast, the high sensitivity of SVD to quantization errors, which arise during a separate quantization step, significantly degrades the reconstruction quality.
@@
As observed in Figure 6(a), contracting the QMF factor bounds from $[-128,127]$ to $[-16,15]$ consistently improves the rate-distortion performance. Generally, narrowing the factor bounds $[\alpha,\beta]$ can potentially lower the entropy, thereby improving the effectiveness of lossless compression in the final stage of our framework and subsequently reducing the bit rate. However, this reduction in entropy comes at the cost of increased reconstruction error, as the feasible set in (3.3) becomes more constrained. This trade-off between entropy and reconstruction quality limits the compression performance of QMF. Therefore, it would be beneficial to moderately expand the factor bounds $[\alpha,\beta]$ while simultaneously controlling the entropy of the elements in the factor matrices. We plan to address this in the future by incorporating an entropy-aware regularization term into the current QMF objective function.
@@
Patchification with an appropriate patch size (e.g., $(8,8)$) helps capture local spatial dependencies and, as confirmed by our results in Figure 6(c), positively impacts the performance of QMF and SVD. However, discontinuities at patch boundaries can introduce blocking artifacts, similar to JPEG compression at very low bit rates (see the building image example in Figure 4). Moreover, while JPEG suffers more from color distortion (e.g., color bleeding and color banding) at low bit rates, QMF and SVD are more affected by blurriness, as observed in the seascape image example in Figure 4. As a potential solution for future work, a deep neural network could be trained to remove these artifacts and then integrated as a post-processing module to further enhance the quality of QMF-compressed images.
6 Conclusion
@@
This work presents a novel lossy image compression method based on quantization-aware matrix factorization (QMF). By representing image data as the product of two smaller factor matrices with elements constrained to bounded integer values, the proposed QMF approach effectively integrates quantization with low-rank approximation. In contrast, traditional compression methods such as JPEG and SVD consider quantization as a separate step, where quantization errors cannot be incorporated into the compression process. The reshaped factor matrices in QMF are compatible with existing lossless compression standards, enhancing the overall flexibility and efficiency of our method. Our proposed iterative algorithm, utilizing a block coordinate descent scheme, has proven to be both efficient and convergent. Experimental results demonstrate that the QMF method significantly outperforms JPEG in terms of PSNR and SIMM at low bit rates and maintains better visual semantic information. This advantage underscores the potential of QMF to set a new standard in lossy image compression, bridging the gap between factorization and quantization.
Acknowledgments
This research received funding from Flanders AI Research Program. Sabine Van Huffel and Pooya Ashtari are affiliated to Leuven.AI - KU Leuven institute for AI, B-3000, Leuven, Belgium.
Appendix A Proof of Theorem 1
@@
$\operatorname*{arg\,min}_{u^{r}_{i}\in\mathbb{Z}_{[\alpha,\beta]}}~{}q_{i}(u_{i}^{r}),\quad\textstyle\text{where}~{}~{}q_{i}(u_{i}^{r})\triangleq\sum_{j=1}^{N}(e_{ij}^{r}-u_{i}^{r}v_{j}^{r})^{2},\quad\forall i\in\{1,\dots,M\}.$ (11)
@@
Since $\bm{u}_{r}^{\star}\triangleq(u^{r^{\star}}_{1},\dots,u^{r^{\star}}_{M})$ is the global optimum of optimization (10), it is evident that
@@
Appendix B Proof of Theorem 2
@@
$\Psi(\bm{U},\bm{V})$ $\triangleq f_{0}(\bm{U},\bm{V})+\sum_{r=1}^{R}f(U_{:r})+\sum_{r=1}^{R}g(V_{:r}),$
$f_{0}(\bm{U},\bm{V})$ $\triangleq\|\bm{X}-\bm{U}\bm{V}^{\rm T}\|_{\rm F}^{2},$
$f(U_{:r})$ $\triangleq\delta_{[a,b]}(U_{:r})+\delta_{\mathbb{Z}}(U_{:r}),$
$g(V_{:r})$ $\triangleq\delta_{[a,b]}(V_{:r})+\delta_{\mathbb{Z}}(V_{:r}),$
@@
The unconstrained optimization problem (14) consists of the sum of a differentiable (smooth) function $f_{0}$ and nonsmooth functions $f$ and $g$. This problem has been extensively studied in the literature under the class of nonconvex nonsmooth minimization problems. In Algorithm 1, the blocks $U_{:r}$ and $V_{:r}$ are updated sequentially following block coordinate descent (BCD) minimization algorithms, also often called Gauss-Seidel updates or alternating optimization [35, 36]. Hence, in this convergence study, we are interested in algorithms that allow BCD-like updates for the nonconvex nonsmooth problem of (14) [37, 38]. Specifically, we focus on the proximal alternating linearized minimization (PALM) algorithm [38], to relate its convergence behavior to that of Algorithm 1. To that end, we show that the updates of Algorithm 1 are related to the updates of PALM on the recast problem of (14), and all the assumptions necessary for the convergence of PALM are satisfied by our problem setting. It is noted that, for the sake of presentation and without loss of generality, in this proof, we assume each of the matrices $\bm{U}$ and $\bm{V}$ has only one column ($R=1$); hence, we only have two blocks in the BCD updates. The iterates in PALM and the presented proof can be trivially extended for more than two blocks.
@@
where the proximal map for an extended proper lower semicontinuous (nonsmooth) function $\varphi:\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ and $\gamma>0$ is defined as $\operatorname{prox}^{\varphi}_{\gamma}(\bm{x})\triangleq\operatorname*{arg\,min}_{\bm{w}\in\mathbb{R}^{n}}\left\{\varphi(\bm{w})+\frac{\gamma}{2}\|\bm{w}-\bm{x}\|^{2}_{2}\right\}$. In (15), $c_{k}>L_{1}(\bm{V}^{k})$ and $d_{k}>L_{2}(\bm{U}^{k+1})$ where $L_{1}>0$, $L_{2}>0$ are local Lipschitz moduli, defined in the following proposition.
The following proposition investigates the necessary assumptions (cf. [38, Asm. 1 and Asm. 2]) for convergence of iterates in (15).
Proposition 1 (Meeting required assumptions).
@@
    $\|\big{(}\nabla_{\bm{U}}f_{0}(\bm{U}_{1},\bm{V}_{1})-\nabla_{\bm{U}}f_{0}(\bm{U}_{2},\bm{V}_{2}),\nabla_{\bm{V}}f_{0}(\bm{U}_{1},\bm{V}_{1})-\nabla_{\bm{V}}f_{0}(\bm{U}_{2},\bm{V}_{2})\big{)}\|\leq M\|(\bm{U}_{1}-\bm{U}_{2},\bm{V}_{1}-\bm{V}_{2})\|,$
@@
By Proposition 1, the optimization problem (14) can be solved by the iterates in (15), due to the following proposition:
Proposition 2 (Global convergence [38]).
With the assumptions in proposition 1 being met by the problem (14), let $((\bm{U}^{k},\bm{V}^{k}))_{k\in\mathbb{N}}$ be a sequence generated by the iterates in (15). Then the sequence converges to a critical point $(\bm{U}^{\star},\bm{V}^{\star})$ of the problem (14), where $0\in\partial\Psi(\bm{U}^{\star},\bm{V}^{\star})$, with $\partial$ as the subdifferential of $\Psi$.
@@
Lemma 1 ($\operatorname{prox}$ implementation).
@@
$\|\bm{W}\|_{[a,b]}^{2}\triangleq\sum_{i,j\mid a\leq\bm{W}_{ij}\leq b}\bm{W}_{ij}^{2},\quad\|\bm{W}\|_{a}^{2}\triangleq\sum_{i,j\mid\bm{W}_{ij}<a}\bm{W}_{ij}^{2},\quad\|\bm{W}\|_{b}^{2}\triangleq\sum_{i,j\mid\bm{W}_{ij}>b}\bm{W}_{ij}^{2}.$
@@
Now that the equivalence of iterates (15) with the simple and closed-form steps in Algorithm 1 is fully established, and the assumptions required for the convergence are verified in proposition 1 to be met by problems (14) and (3.3), proposition 2 can be trivially invoked to establish the convergence of Algorithm 1 to a locally optimal point of problem (3.3).